OpenStudy (littlet):

What is the equation of a circle with center at (3, -5) and a radius of 7?

OpenStudy (anonymous):

first, do you know the equation of a circle?

OpenStudy (littlet):

If i do i don't remember it.

OpenStudy (anonymous):

(x – h)^2 + (y – k)^2 = r^2

OpenStudy (littlet):

Alright

OpenStudy (anonymous):

(h,k) is the coordinates for the center of the circle

OpenStudy (littlet):

Ok.

OpenStudy (anonymous):

(h,k) could be viewed as (x,y) if you want so just substitute 3 for h and -5 for k and r for 7

OpenStudy (anonymous):

(x – 3)^2 + (y + 5)^2 = 7^2

OpenStudy (littlet):

Alright.

OpenStudy (anonymous):

(x – 3)^2 + (y + 5)^2 = 7^2 =49

OpenStudy (anonymous):

tada! you're done

OpenStudy (littlet):

Thank you so much!